LjTheBaller16:

x y 3 10 20 In the table, the relation (x, y) is not a function if the missing value of x is .

LjTheBaller16:

the 20 is in the right spot

darkknight:

linear function?

darkknight:

I meant is this function proportional

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @darkknight I meant is this function proportional \(\color{#0cbb34}{\text{End of Quote}}\) wait wym by dat

Vocaloid:

Major hint: a relation is *not* a function if one input leads to multiple outputs at once So which x-value would make the relation *not* a function?

jhonyy9:



darkknight:

we are not quite understanding the table, mind uploading it (attach file button)

jhonyy9:

@darkknight re-read pls the text - there say about missing value of x

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 Created with Raphaëlxy31020?Reply Using Drawing \(\color{#0cbb34}{\text{End of Quote}}\) yea dats how it look

darkknight:

ah my b \(\color{#0cbb34}{\text{Originally Posted by}}\) Vocaloid Major hint: a relation is *not* a function if one input leads to multiple outputs at once So which x-value would make the relation *not* a function? \(\color{#0cbb34}{\text{End of Quote}}\) she said this hint perfectly

jhonyy9:

so i think you can use the rule of 3 simple and write in this way when x = 3 result y =10 so for x = ,,a" result y = 20 cross multiplie and calcule ,,a" a = 20*3/10 = 60/10 = 6 for x = 6 will be y = 20 hope helped easy understandably

jhonyy9:

@Vocaloid

Vocaloid:

Keep in mind what the question is asking It’s not asking us to calculate a missing value that would complete a linear relationship It’s asking for an x-value that would make the relation *not* a function

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid Keep in mind what the question is asking It’s not asking us to calculate a missing value that would complete a linear relationship It’s asking for an x-value that would make the relation *not* a function \(\color{#0cbb34}{\text{End of Quote}}\) they have a good point

Vocaloid:

This is hard to explain without giving it away, but think about what would happen if we tried assigning x = 3 to another value in addition to y = 10

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid This is hard to explain without giving it away, but think about what would happen if we tried assigning x = 3 to another value in addition to y = 10 \(\color{#0cbb34}{\text{End of Quote}}\) you tryna say i need to multiple

Vocaloid:

You don’t need to calculate anything for this problem Notice how x = 3 is assigned to y = 10 If we also assign x = 3 to y = 20, it’s not a function anymore, because we assigned multiple outputs (y = 10 and y = 20) to the same input (x = 3)

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid You don’t need to calculate anything for this problem Notice how x = 3 is assigned to y = 10 If we also assign x = 3 to y = 20, it’s not a function anymore, because we assigned multiple outputs (y = 10 and y = 20) to the same input (x = 3) \(\color{#0cbb34}{\text{End of Quote}}\) im confused

Vocaloid:

Look at the first row of the table x = 3 matches with y = 10 Now look at the second row of the table x = ? And y = 20 If we let “?” be 3 That means x = 3 goes with y = 10 AND y = 20 at the same time This violates our rule about functions (each x value can only match with one y value at a time)

Vocaloid:

Since the question is asking what missing x-value will make the relation *not* a function, that means our missing value is x = 3

LjTheBaller16:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Vocaloid Since the question is asking what missing x-value will make the relation *not* a function, that means our missing value is x = 3 \(\color{#0cbb34}{\text{End of Quote}}\) oh okay i gotchu

jhonyy9:

ty. @Vocaloid this is clearly now